Please help me with this problem: 4+|3-(x/3)|>=9
I get x<=-6 and x>=48 my professor says that the 48 should be 24 and i just don't understand why!

Answer by Alan3354(69443)

4+|3-(x/3)|>=9
|3-(x/3)| = 5
--------------
3-(x/3) = 5
x/3 = -2
x = -6
-----------
(x/3)-3 = 5
x/3 = 8
x = 24
--------
How did you get 48?
-------------------
I get x<=-6 and x>=48 my professor says that the 48 should be 24 and i just don't understand why